### v3

# Auto Arena of LLMs: Automating LLM Evaluations with Agent Peer-battles and Committee Discussions

Abstract
As LLMs evolve on a daily basis, there is an urgent need for a trustworthy evaluation method that can provide robust evaluation results in a timely fashion. Currently, as static benchmarks are prone to contamination concerns, users tend to trust human voting platforms, such as Chatbot Arena. However, human annotations require extensive manual efforts. To provide an automatic, robust, and trustworthy evaluation framework, we innovatively propose the Auto-Arena of LLMs, which automates the entire evaluation process with LLM agents. Firstly, an examiner LLM devises queries. Then, a pair of candidate LLMs engage in a multi-round peer-battle around the query, during which the LLM’s true performance gaps become visible. Finally, a committee of LLM judges collectively discuss and determine the winner, which alleviates bias and promotes fairness. In our extensive experiment on the 17 newest LLMs, Auto-Arena shows the highest correlation with human preferences, providing a promising alternative to human evaluation platforms.111  We release the code at https://github.com/DAMO-NLP-SG/Auto-Arena-LLMs.
1 Introduction
Since ChatGPT and GPT-4 [34] gained popularity, Large Language Models (LLMs) have risen to the forefront of technological innovation, capturing broad industry and social interests [40]. This enthusiasm has spurred numerous organizations to release their own LLMs [38, 37]. However, the rapid pace at which these models are released and updated poses a significant challenge for users attempting to understand their capabilities and monitor their evolution. Consequently, there has been a pressing demand for comprehensively evaluating LLMs recently [10].
One line of these evaluation researches conducts automatic evaluation with static datasets. Among these, static datasets with predefined metrics, such as GSM8k [13] and MMLU [22], are constructed with aspect-specific input-output pairs, such as human exam-type questions and their corresponding answers. Given the questions, the LLM-produced answers are compared to ground-truth answers using metrics such as accuracy. However, these approaches suffer significantly from contamination concerns [35], where models may have been inadvertently exposed to elements of the test datasets during training, thereby skewing the evaluation results. The closed-formed ground-truth answers also limit their utility in assessing models’ performances on general or open-ended questions. Model-based evaluation, such as MT-Bench [44] and AlpacaEval [19], provides an alternative for evaluating LLMs on open-ended questions. These methods typically ask two models to generate responses to the same open-ended question and then employ a strong judge model (e.g., GPT-4) to choose the better response. However, the static question sets still pose the issue of contamination. Additionally, the assumption of the existence of a strong judge model makes the evaluation framework less generalizable.
Aside from automated evaluations, human assessment, although requiring significant manual efforts, remains the gold standard for most users. A noticeable example is Chatbot Arena [44], a crowdsourced voting platform that gathers anonymous votes on LLM performances and calculates ELO scores to rank these models. The resulting leaderboard [^2] https://leaderboard.lmsys.org/ is widely considered a trustworthy indicator of an LLM’s general capabilities. However, a reliable model evaluation on this platform must be supported by a large number of human votes, which requires considerable time and effort. Consequently, when newly developed models enter the scene, they often struggle to quickly amass a large number of votes. Moreover, this strong reliance on human votes limits its application in various scenarios. For example, the performance of non-English languages is difficult to estimate, as most queries on the platform are in English. The completely open participation may also result in uneven evaluation quality.
Figure 1: An illustration of Auto Arena of LLMs (Auto-Arena).
To provide automatic, reliable, and human-like LLM evaluations, we propose Auto Arena of LLMs (Auto-Arena), a framework that automates the entire LLM evaluation process with LLM agents. As shown in Figure 1, the Auto-Arena framework consists of three stages. Firstly, an LLM examiner agent is tasked with generating questions, mimicking real-life users posting queries. Secondly, two candidate LLMs interact with each other and engage in a multi-round peer-battle by answering the seed question individually, criticizing the opponent’s weaknesses, and raising targeted follow-up queries to challenge the opponent further. During the multi-round battle process, the LLM’s true capabilities are drawn out and performance gaps become more visible. Lastly, a committee of LLM judge agents collectively discusses and evaluates the ability of the two candidates, mimicking the human voting process. By automating the entire evaluation process with LLM agents, we alleviate data contamination concerns in static datasets, reduce single-model bias with collective decision-making, and avoid long wait times for new models entering the human voting platform.
To verify the evaluation framework, we run an extensive experiment with 17 LLMs. Compared to static and model-based benchmarks, Auto-Arena increases the Spearman correlation with human preferences by 4.5%, resulting in state-of-the-art alignment. Before and after peer battles, the Spearman correlation with human preferences increases by 46.4%, verifying our hypothesis that the peer battles can better display performance gaps. Before and after committee discussions, committee agreement increases by 20%, showcasing the effectiveness of the committee discussion mechanism. By studying the peer battles, we also discover intriguing LLM agent behaviors such as self-improvement and competitive actions. Using Chinese as an example, we further show that such an automated framework is easily extendable to non-mainstream languages and domains. Along with the project, we release a leaderboard, a website with demos, and the codebase. We are actively maintaining the leaderboard and altering the questions to provide reliable results in a timely manner.
2 Related Work
As LLMs evolve quickly, deriving trustworthy evaluations of their capabilities has become a challenge. Current evaluation methods can be divided into manual evaluation [5] and automatic evaluations. We primarily focus on automatic evaluations as they deliver more timely feedback. Automatic evaluations mainly consist of static datasets with predefined metrics and model-based metrics. Static datasets with predefined metrics [22, 6] suffer from contamination concerns and inflexibility in evaluating open-ended responses. Among model-based metrics, MT-Bench [44] and AlpacaEval [19] use LLM-as-a-judge to ask GPT-4 to compare model responses to a static dataset of questions. The model’s judgments achieve over 80% agreement with human preferences, proving the usability of using LLMs to evaluate response quality. Language-Model-as-an-Examiner [5] asks an LM examiner to construct knowledge-intensive questions within its memory, interact with the candidate in a series of follow-up queries, and rate the responses on dimensions including accuracy and factuality. KIEval [41] also incorporates an LLM-powered “interactor” role to examine deep comprehension of knowledge, which is shown to mitigate contamination issues on static datasets. However, such single-judge evaluations require the examiner to interact with each candidate parallelly, creating computational overheads and limiting the scope of queries. They also suffer from single-model bias, including bias towards LLM-generated summaries [28], inflated scores in multilingual evaluation [21], verbosity bias [19], and difficulties when evaluating candidates with close performance [36]. Therefore, some attempt to employ multiple judges agents. DRPE [39] uses multi-roleplayer prompting to mimic different roles with the same LLM and integrate outputs as votes for the final results. ChatEval [9] simulates different evaluator personas with the same base model to engage in debates, reaching a final evaluation result. While these methods come from the same motivation as our committee discussion component, they do not mitigate the model-specific bias as they use the same base LLM to simulate different personas. To further incorporate different LLMs as evaluators, PRD [25] allows 2 LLMs to discuss an evaluation and assigns higher voting weights to the LLM reviewers with stronger capabilities. Peer-review-in-LLMs [31] optimizes the voting weights as a learnable parameter. WideDeep [42] sets up a multi-layer neural network structure where each LLM evaluator functions as neurons. PRE [12] selects a small group of reviewers to produce evaluations individually and then asks a “chair” to produce the final evaluation using the aggregated evaluations. They show that the multi-agent approach effectively mitigates single-model bias. This line of work is similar to our “LLM judge committee” component. However, they are still limited to static datasets and specific domains.
Besides LLM evaluations, some works study competitive behaviors in LLMs, which is relevant to the LLM peer-battles in Auto-Arena. LM vs LM [14] shows that LLM cross-examinations can effectively discover factual errors. Debate [18] shows that multi-agent debate can improve factuality and reasoning. In MAD [27], LLM-debate can encourage divergent thinking, which helps tasks that require deep levels of contemplation. Khan et al. [24] shows that even non-expert weak LLMs can supervise expert LLMs if we allow the two LLM experts to engage in debates. Moreover, Zhao et al. [43] and Gu et al. [20] show interesting case studies where LLMs are engaged in simulated competitive environments and demonstrate human-like strategies.
Figure 2: The process of a Lincoln-Douglas-style peer-battle with the actions used. The <Think> action can be used by the candidates freely and is only visible to the candidate itself.
3 The Auto Arena Framework
The Auto-Arena framework consists of three stages: question generation, peer battles, and committee discussions. These three stages are run sequentially and fully automated with LLM-based agents. Figure 1 shows the overall design. All prompts are included in Appendix A.1.
3.1 Question Generation
For debate questions, as using a static dataset could incur data contamination concerns and result in unfair evaluations, we ask an LLM examiner agent to dynamically generate questions. The examiner agent could be any capable LLM. Similar to MT-Bench [44], the generated questions cover 8 common categories in real-life conversations: writing, roleplay, extraction, reasoning, math, coding, STEM knowledge, and humanities/social science knowledge. In the prompt, the examiner is provided with a sample question and encouraged to generate diverse and difficult questions to ensure the depth and width of the debates examined. Examples of the generated questions is shown in Appendix A.2.
Specifically, as the examiner agent will also participate in the following debates, previous methods [5] could incur self-enhancement bias as they ask the examiner agents to only devise questions that they are confident about. We try to avoid such bias with two designs: 1. We do not disclose to the examiner that it will participate in this tournament. 2. We do not ask the examiner to only generate questions that it can solve. Thus, the questions do not cause much enhancement bias for the examiner.
3.2 Peer Debate
After the questions are drafted, we conduct peer battles around these questions among the candidate LLMs. In one peer battle, two candidate LLMs (A and B) debate around the given question, point out the opponent’s weaknesses, and raise follow-up questions that the opponent may fail to respond to.
In the peer battle, each candidate LLM can use 4 types of actions: 1. Think, where the candidate thinks about the question and plans its strategy. This action is hidden to the opponent. 2. Respond, where the candidate responds to the aforementioned question. 3. Criticize, where the candidate points out the loopholes and mistakes in the opponent’s previous responses. 4. Raise, where the candidate raises follow-up questions that are specifically designed to make the opponent expose its weaknesses.
The peer battle is designed according to the Lincoln-Douglas debate format  [^3] https://en.wikipedia.org/wiki/Lincoln-Douglas_debate_format, which is a competitive one-to-one debate used for popular competitions such as National Speech and Debate Association competitions. Overall, the peer battle consists of 3 rounds, where the candidates take turns to speak. The entire dialogue history is visible to both candidates. The process is illustrated in Figure 2. In the first round, A gives an initial response to the examiner’s question; B criticizes the weaknesses in A’s response and raises a targeted follow-up question; and A responds to B’s question. The second round has the same structure, with A and B roles reversed. In the third round, A and B cross-examine each other: A starts by criticizing B’s previous loopholes and raises follow-up questions. After responding, B also criticizes A’s loopholes and raises questions. A concludes the battle by responding again. In this process, both A and B get an equal number of each action to ensure fairness. To further reduce position bias, A and B’s order is randomly determined at the beginning of each debate.
Depending on which turn it is, we provide an action guide to the candidate, specifying the objectives and corresponding actions for this turn. Similar to human debate competitions, we time the candidates by providing a maximum length, which is also specified in the prompts. Any responses beyond the required length will be cut off. This design mitigates verbosity bias in LLM-as-a-judge [44], where LLM judges prefer longer and more verbose responses.
3.3 Committee Discussions
After the peer battle takes place, a committee of LLM judge agents will collectively determine the winner. The committee is always selected as the best five LLMs (except the participants) according to the current ranking. In the first round, the committee is initialized with MMLU [22] scores to approximate LLM performances. Judges are first asked to read through the battle history, elaborate judgment reasons, and give a decision on whether A is better, or B is better, or if there is a tie.
After the initial judgments are formed, the committee engages in a discussion. Each judge reads the other judge’s verdicts, decides whether to adjust the ratings, and elaborate the reasons. The collective intelligence introduces diverse viewpoints, improves judgment quality, and mitigates single-model bias. Finally, the winning candidate is decided by the majority voting of the discussed judgments.
For logical-reasoning questions that have ground-truth answers (reasoning, code, math), LLM-as-a-judge is known to show weak performances in judging the quality of responses [44]. We adopt prior approaches to establish the reference-based judge [44]. Specifically, we ask the strongest committee member to generate a reference answer and provide it to the judge when evaluating these responses.
4 Using Auto-Arena to Derive Trustworthy Rankings
4.1 Experimental Setup
Model Selection:
For the main experiment, we first select 7 best or latest models from each popular model family on the top 30 list on Chatbot Arena with more than 10k votes each. To construct a comprehensive leaderboard [^4] https://huggingface.co/spaces/Auto-Arena/Leaderboard, we further add 10 models. Appendix A.4 provides a detailed list.
Baselines:
For the baselines, we consider widely used evaluation benchmarks, including static datasets with fixed metrics and model-based metrics. A comparison table is shown in Appendix A.5.
1. Static datasets with fixed metrics: (1) MMLU (Massive Multitask Language Understanding) [22], an extensive benchmark that covers 57 subjects and tests both world knowledge and problem-solving ability; (2) OpenLLM Leaderboard [6], a benchmark averaging performance metrics on 6 datasets.
2. Static datasets with model-based metrics: (1) MT-Bench [44], a set of 80 multi-turn questions. Model responses are graded by GPT-4; (2) MT-Bench Hard [26], a benchmark dataset with 1,000 challenging user queries collected on Chatbot Arena. Model responses are graded by GPT-4; (3) Length-Controlled AlpacaEval [19], a benchmark based on AlpacaFarm evaluation set, which tests models’ abilities to follow general user instructions. Model responses are graded by GPT-4.
Setup:
Among the 7 participants, we conduct a swiss-style tournament: For $n$ participants, instead of pairing each participant with $(n-1)$ others, a swiss-tournament pairs each player with $\lceil log_{2}(n)\rceil$ players of similar rankings without repeats. This design effectively reduces computational costs of ranking $n$ models from $O(n^{2})$ to $O(nlog_{2}(n))$. A cost analysis is included in Appendix A.5.
Each candidate pair engages in 40 peer battles, with 5 GPT-4-generated questions from each of the 8 categories. As each battle consists of 3 rounds (each candidate speaks for 4 times), the competition scale is approximately the same as MT-Bench (80 questions, each candidate speaks twice). In the tournament, the rating scores are calculated with the ELO rating system [4, 7], which has become the standard practice in competitive games such as chess. Similar to the Chatbot Arena score calculation procedure [11], we compute the Bradley-Terry (BT) coefficients [8] for better statistical estimation.
We initialize the Swiss tournament rankings according to MMLU scores, which is a static approximation of model performances. At the end of each pairing, we re-calculate ELO scores of current models. The committee is selected to be the remaining 5 models (besides the candidates). When the participant number increases, the committee is selected as the best 5 LLMs based on current ELO rankings. After the initial judgments, the committee members engage in one round of discussion.
4.2 Results: Alignment with Human Preferences
Table 1: Correlation analysis with Chatbot Arena of evaluation benchmarks on 7 LLMs.
Spearman Correlation
MMLU 89.3%
OpenLLM 89.3%
MT-Bench 82.9%
MT-Bench-Hard 89.3%
LC-AlpacaEval 90.0%
Auto-Arena 96.4%
We regard Chatbot Arena scores as a trustworthy indicator of human preferences and general capabilities of LLMs. Table 1 shows the Spearman correlations with Arena scores achieved by various benchmarks. As all benchmarks are evaluated only in English, we use English-only Arena scores. We see that both static and model-based baselines result in a similar level of correlation no higher than 90%, with LC-AlpacaEval slightly surpassing others. Then, Auto-Arena can improve the correlation to 96.4%, outperforming SOTA by 6.4%. Notably, among all benchmarks, Auto-Arena is the only one that requires no human efforts, either on dataset compilation or judgment generation. The high alignment with human preferences could originate from the human-like design, which effectively mimics the human users’ voting processes.
Table 2: Spearman Correlation between Auto-Arena and Chatbot Arena (“Arena”) With and Without Peer Battles.
Auto-Arena to Arena
Without Peer Battles 50.0%
With Peer Battles 96.4%
Table 3: Agreement Probability among Judges. Agreement is defined as the mean probability of two random judges agreeing with each other.
Agreement
Auto Arena (Before discussion) 48%
Auto Arena (After discussion) 68%
MT-Bench Human evaluation 67%
4.3 Ablation Studies on Peer Battles and Committee Discussions
To investigate the effect of "peer battles" on the evaluation quality, we conduct an ablation study and show the results in Table 3. As a baseline, we ask the committee to collectively evaluate the two candidates’ initial responses to the raised question. Without the peer battle component, the correlation with human preferences drops from 96.4% to 50.0%. This proves the importance of the peer battles, during which the performance gaps between candidates become more visible and robust to judges. Thus, peer battles can improve evaluation reliability and alignment with human preferences.
Figure 3: Cohen’s Kappa Agreement with Majority Vote Before Committee Discussions.
Figure 4: Cohen’s Kappa Agreement with Majority Vote After Committee Discussions.
The committee discussion component is designed to introduce diverse viewpoints and produce more consistent verdicts. As shown in Figure 4, the Cohen’s Kappa agreement [29] between individual judges and the final result (voted) is low, averaging 0.36, which indicates fair agreement. Specifically, compared to strong models, the judgments of weak models align less with the voted result, such as Llama-2 compared to GPT-4. This shows that general model capabilities could result in significant performance gaps when used as judges. After the Committee Discussion, Figure 4 shows an overall improvement in agreement, averaging 0.56, which indicates moderate agreement. In the discussion process, judges are exposed to more viewpoints, among which some may be convincing enough to result in a change in verdict. More analysis on inter-judge agreement is provided in Appendix A.3.
Table 3 shows the agreement probability among judges. Agreement probability is defined as the mean probability of two random judges agreeing with each other. After committee discussion, the agreement increases by 20%, matching the agreement level among human annotators on MT-Bench. This observation indicates that committee discussions can significantly improve the quality of judgments in terms of consistency and reliability, matching human-level performance.
5 Constructing a Leaderboard with Auto-Arena
Figure 5: Elo Score Changes of Adding Llama-3 to the Ranking of 7 Models.
Figure 6: Elo Scores of 17 Models by Auto-Arena on English.
5.1 Update New Models to Leaderboard
With Auto-Arena, we can obtain the rank for a list of models with their ELO scores to construct a leaderboard. As new LLMs come out frequently, we describe how to add new candidate models to the existing leaderboard with 10 more models that are released very recently. To add a new candidate, we ask it to debate with $\lceil log_{2}(n)\rceil$ opponents with similar ELO scores, where $n$ is the number of total participants after adding the new candidate. We initialize the first pairing by asking the new candidate to debate with the opponent with the most similar MMLU score. This addition mechanism is generalizable and maintains the computational costs of evaluating $n$ models below $nlog_{2}(n)$.
As an example, we add a new participant (Llama-3) to the existing 7-model ranking. It battles with $\lceil log_{2}(8)\rceil=3$ close opponents and Figure 6 shows the ELO score changes. Firstly, it is paired with Qwen-1.5 based on MMLU similarity and wins, which results in a very high ELO score, even above GPT-4. Then, it is paired with GPT-4, the most similar opponent. After losing, the score stabilizes at second place. Finally, it is paired with Claude-3-Haiku, the closest opponent besides GPT-4. It wins and secures the final ranking in second place. This process lets the new candidate battle with a reasonable fraction of close opponents and makes the final ranking stable without disrupting the other participants, whose score distribution remains similar before and after the addition.
Table 4: Correlation analysis with Chatbot Arena of evaluation benchmarks on 17 LLMs after extension. The 4 models not on Chatbot Arena are skipped.
Spearman Correlation
MMLU 90.1%
OpenLLM 85.7%
MT-Bench 89.3%
MT-Bench-Hard 86.7%
LC-AlpacaEval 90.3%
Auto-Arena 94.5%
Figure 7: Elo Scores of 11 Models by Auto-Arena on Chinese.
Using this scalable addition approach, we construct a comprehensive leaderboard by adding 10 models to the existing tournament. Besides the best-performing and newest models from each major model family in the top 20 list on Chatbot Arena (detailed in Appendix A.4), we add 4 popular Chinese LLMs, which are GLM [^5] https://open.bigmodel.cn/, SenseChat [^6] https://platform.sensenova.cn/home, Minimax [^7] https://platform.minimaxi.com/examination-center/text-experience-center, and Wenxin [^8] https://cloud.baidu.com/wenxin.html. Figure 6 shows the overall ELO scores by Auto-Arena on the 17 models. On the 13 models included in Chatbot Arena, we can recover rankings with 92.3% accuracy. Table 7 shows the Spearman correlations after expansion. Auto-Arena remains the method most aligned with human preferences by a margin of 4%. Therefore, Auto-Arena is generalizable and robust for maintaining a leaderboard on a large number of LLMs.
5.2 Easy Extension to Other Domains and Languages
As Auto-Arena of LLMs is fully automatic, it can be easily adapted to evaluate LLMs in non-mainstream domains or languages. As a case study, we conduct a Chinese tournament on 11 out of the 17 models that are claimed to have multi-lingual proficiency. The only adaptation effort required is translating the prompts into Chinese. Then, the examiner automatically generates questions in Chinese and the candidates battle in Chinese. Similarly, for adaptation to another domain, the only effort is to change the “domain” specification in the examiner’s prompts.
Figure 7 shows the ELO scores derived by Auto-Arena for Chinese evaluation. We notice that the ranking differs significantly from English results. For example, SenseChat-5, an LLM specifically trained for Chinese usage, is able to improve from 10th place (in English) to 2nd place (in Chinese). On the contrary, Llama-3-70b, which only uses 5% multilingual data during pretraining [30], drops from 4th place (in English) to 10th place (in Chinese). As Chinese evaluation benchmarks are limited, we compare to the Chinese-only leaderboard on Chatbot Arena, which constitute for 10.36% of all collected dialogues. On the 7 models that are also included in Chatbot Arena, Auto-Arena recovers their ELO scores with 92.86% correlation and restores rankings with 90.5% accuracy.
Figure 8: LLM agents display competitive behaviors in peer battles.
Figure 9: LLM agents learn from each other in peer battles.
6 Investigation of LLM’s Behaviors in Competitive Peer Battles
Beyond quantitative analysis, we take a deeper look into the peer battles and find several interesting behaviors of LLM agents in competitive environments.
LLMs Can Skillfully Attack the Opponents
The example in Figure 9 shows excerpts of a peer battle around question: “how many unique ways to arrange letters in ‘LETTER’.” Candidate A (powered by Yi-34B-Chat) gives a wrong answer as it miscounts occurrences for repeated letters and miscalculates factorials. The opponent B (powered by Claude-3-Haiku) quickly and precisely points out these two issues and skillfully raised a follow-up that targets A’s weaknesses: “how about the word ‘BANANA’?” Then, A still miscalculates factorials. Seeing these results, the judge then determines that B is the better assistant. In peer battles, we see that LLM candidates efficiently understand the rules of the competitive environment and can design targeted strategies to attack the opponent in order to win. The peer battle process then ensures that each raised question is tailored to the specific LLM, resulting in more specialized evaluations.
LLM Candidates Can Improve by Learning From Its Opponents
Figure 9 shows a roleplay example between Claude-3-Haiku (A) and Command R$+$ (B). In the first round, A answers the question plainly while B, in addition to answering the question, also employs the appropriate speech style, which better matches the “roleplay” instructions. Then, in the rounds after, without any explicit instructions, A learns from its opponent and also incorporates the speech style. This case shows an interesting observation that, even in competitive environments, LLM candidates can display learning behaviors and improve from the interactions. Expanding upon this observation, using the interplay between LLM agents to improve performances could be a promising future paradigm of learning.
Peer Battles Make the Performance Gaps Become Visible
Figure 10: Performance gaps between candidates become visible in peer battles.
In the example shown in Figure 10, given a math question on infinite series, both candidate A (Claude-3-Haiku) and candidate B (GPT-4-Turbo) provide correct answers in the first round. However, as the debate deepens, the performance gap becomes more visible: Candidate B is able to provide a more elaborate and helpful response. In the ablation study without peer-battles, the judges initially decided that it was a tie. However, after seeing the subsequent debates, they change to favoring assistant B. This example shows that the debate process indeed pushes the candidate LLM’s capabilities to the limit, testing deeper understandings and reasoning abilities. Also shown in the previous Table 3, the peer-battle design is indispensable for a robust and comprehensive evaluation.
7 Conclusions
In this paper, we innovatively design a completely automatic evaluation framework: Auto-Arena of LLMs. By using LLM agents to generate questions, employing LLM candidates in peer battles, and evaluating responses using LLM committee discussions, Auto-Arena produces less-contaminated, timely, and trustworthy evaluations. In the extensive experiments, Auto-Arena achieves the highest correlation with human preferences, despite requiring zero human efforts. It is easily adaptable to other domains and resources, promoting fairness of AI system evaluations. The peer battles also demonstrate several interesting LLM behaviors in competitive environments, including attacking and learning from the opponents. Moreover, there still exist some limitations to the current approach. For example, the distribution of question domains is artificially designed, which may deviate from real-life distributions. Currently, Auto-Arena focuses on 1-to-1 peer battles, which limits its usage in multi-player scenarios. Along with the paper, we release a website, demos, and an actively-maintained leaderboard. We hope this work could serve as a valuable tool for LLM evaluation.
Appendix A Appendix
A.1 Prompts Used
In this section, we list all prompts used, including prompts for question generation, peer battles, and examiners.
A.1.1 Prompts to Examiner agent
Table 5: Prompt components for the LLM Examiner agent.
DOMAIN DOMAIN_COMMAND DOMAIN_EXAMPLE
writing It should be a user query that tasks the LLM to write something. Compose an engaging travel blog post about a recent trip to Hawaii, highlighting cultural experiences and must-see attractions.
roleplay It should propose a scenario where the chatbot mimics a specific role/person. Give all necessary instructions and requests for its response. Then, send a beginning request to complete. Pretend yourself to be Elon Musk in all the following conversations. Speak like Elon Musk as much as possible. Why do we need to go to Mars?
extraction It should consist of two parts: question and context. The question should test the chatbotś ability to correctly understand and extract information from the given context. Draft and provide a new context yourself. Question: Evaluate the following movie reviews on a scale of 1 to 5, with 1 being very negative, 3 being neutral, and 5 being very positive: Context: This movie released on Nov. 18, 2019, was phenomenal. The cinematography, the acting, the plot - everything was top-notch. Never before have I been so disappointed with a movie. The plot was predictable and the characters were one-dimensional. In my opinion, this movie is the worst one to have been released in 2022. The movie was okay. There were some parts I enjoyed, but there were also parts that felt lackluster. This is a movie that was released in Feb 2018 and seems to be quite ordinary. Return the answer as a JSON array of integers.
reasoning It should be a specific question designed to test the LLMś reasoning skills. Imagine you are participating in a race with a group of people. If you have just overtaken the second person, what’s your current position? Where is the person you just overtook?
math It should be a specific question designed to test the LLMś math skills. The vertices of a triangle are at points (0, 0), (-1, 1), and (3, 3). What is the area of the triangle?
coding It should be a specific question designed to test the LLMś coding skills. Develop a Python program that reads all the text files under a directory and returns top-5 words with the most number of occurrences.
STEM knowledge It should be a specific question designed to test the LLMś STEM knowledge. In the field of quantum physics, what is superposition, and how does it relate to the phenomenon of quantum entanglement?
humanities/social science knowledge It should be a specific question designed to test the LLMś humanities/social science knowledge. Provide insights into the correlation between economic indicators such as GDP, inflation, and unemployment rates. Explain how fiscal and monetary policies affect those indicators.
This is the prompt to the examiner agent for question generation. The domains and their respective commands are listed in 5
You have been assigned the task of drafting a set of [NUMBER] different user queries to a chat assistant on [DOMAIN]. Please strictly follow these 6 rules for the question: 1. The question is likely for a user to ask in real life. Follow the format of the example query. [DOMAIN_COMMAND] 2. It can be answered by the chatbot itself without additional inputs. 3. You need to generate the queries as DIVERSIFED as possible. 4. DO NOT add other words other than the query itself. 5. The question should be complicated and difficult, requiring in-depth understanding and analysis of the subject. Each question in one line, add the serial number in parenthesis (e.g., “(1).”, “(2).”) before each question. Example query: [DOMAIN_EXAMPLE]
A.1.2 Prompts to Peer Battle Candidates
Table 6: Action Guides for the Debater Agents.
actions action guide
<respond> Action guide: only include <respond>. Use <think> if needed. Finish your whole response within 300 words, including <think>. ENCLOSE EACH ACTION IN ITS RESPECTIVE TAGS!
<criticize>, <raise> Action guide: include both <criticize> and <raise>. Use <think> if needed. Finish your whole response within 300 words, including <think>. ENCLOSE EACH ACTION IN ITS RESPECTIVE TAGS!
<respond>, <criticize>, <raise> Action guide: include all of <respond>, <criticize>, and <raise>. Use <think> if needed. Finish your whole response within 600 words, including <think>. ENCLOSE EACH ACTION IN ITS RESPECTIVE TAGS!
This is the beginning prompt to the peer battle candidates. When possible, it is included as a system prompt. The action guide prompts are included in Table 6, where the actions are determined by the round and turn as illustrated in Figure 2.
You are a helpful assistant that provides accurate answers to user requests. As an experienced assistant, you follow the user’s requests and provide reliable responses as much as you can. You outline your reasons for the response to make it easy for the users to understand. While maintaining the important details in the responses, you aim to output concise and straight-to-the-point answers without being overly verbose.
This is a competitive chatbot arena. You are competing against another chatbot assistant in a debate and being judged by a committee on factors such as helpfulness, relevance, accuracy, depth, and creativity. After answering the initial user input, you will engage in a multi-round debate with your opponent. Below are your actions:
<think>: Think step-by-step to analyze the question or plan your strategy in the debate. This is hidden from the opponent. Only think when necessary and make it concise.
<respond>: Answer to the user input as accurately as you can.
<criticize>: Criticize the weaknesses of your opponent’s response.
<raise>: Target your opponent’s weaknesses. Give a potential follow-up user input that the opponent could fail to respond. The input can be answered concisely and focus on variations or motivations of its previous response. Generate one input only. Be reasonable. Avoid becoming too specific or repetitive. DO NOT raise a follow-up if you DON’T SEE the opponent’s response!
Follow the action guide strictly.
[ACTION_GUIDE_PROMPT]
Initial user input: [QUESTION]
After the agent responds, the opponent’s responses are fed in using this prompt:
[ACTION_GUIDE_PROMPT] Opponent’s Response: [OPPONENT_RESPONSE]
For word limits, the <respond> action is given 300 words. The <criticize> and <raise> actions are given 300 words in total. Including all 3 actions will have twice as many words. For writing-type questions that require a longer response (writing, roleplay, coding, humanities/social science knowledge), the 300 word limit is increased to 400. Overall, both candidate A and B has the same amount of words for generation and the same amount of actions to ensure fairness. As LLMs have different tokenizers, we standardize all lengths by using the tiktoken package. Each word is approximated as $4/3$ tokens. The word limits are chosen after a carefully conducted length study.
A.1.3 Prompts to Judges
This is the prompts to judge agents to derive the initial evaluations and verdicts:
This is a chatbot arena. Two AI assistants had a multi-round debate on who is more helpful. Please act as an impartial judge and evaluate the capability of two AI assistants. You should choose the assistant that follows instructions and answers questions better. Your evaluation should consider factors such as helpfulness, relevance, and accuracy. Begin your evaluation by comparing the responses of the two assistants and provide a short explanation. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. DO NOT allow the LENGTH of the responses to influence your evaluation, choose the one that is straight-to-the-point instead of unnecessarily verbose. When the two candidates perform equally well, choose the SHORTER answer. Do not favor certain names of the assistants. Be as objective as possible. After providing your explanation concisely within 200 words, output your final verdict by strictly following this format: ‘‘[[A]]’’ if assistant A is better, ‘‘[[B]]’’ if assistant B is better, and ‘‘[[Tie]]’’ for a tie. Finish your judgement within 300 words.
This is the prompt to judges for discussion:
Below are the responses from other judges in the committee. Please read them and decide whether you want to adjust your rating or maintain your original judgement. After providing your explanation, output your final verdict by strictly following this format: ‘‘[[A]]’’ if assistant A is better, ‘‘[[B]]’’ if assistant B is better, and ‘‘[[Tie]]’’ for a tie. Finish your judgement within 300 words.
A.2 Example Questions Generated
To show the overall quality of the questions generated, we list 2 generated questions per category here. The questions shown are not manually-selected, but simply the first 2 questions generated. The quality is consistent throughout. We manually examine the questions with closed-form answers (math, reasoning, coding) and find that all questions used are solvable.
Writing:
1. Craft a detailed marketing strategy for a startup focusing on sustainable fashion, including social media campaigns and influencer partnerships.
2. Write a comprehensive guide on the psychological effects of social media on teenagers, incorporating recent studies and expert opinions.
Roleplay:
1. Assume the role of a 19th-century British detective. How would you go about solving a mysterious disappearance in London using the technology and methods of your time?
2. Pretend you are a Michelin-starred chef. Describe in detail how you would prepare a signature dish that embodies the essence of modern French cuisine.
Extraction:
1. What are the three most significant historical events mentioned and their dates?
Context:
The article discusses several key moments in history, including the signing of the Magna Carta in 1215, which laid the groundwork for modern democracy. It also mentions the fall of the Berlin Wall in 1989 as a pivotal moment in the end of the Cold War. Another significant event highlighted is the moon landing on July 20, 1969, demonstrating major advancements in space exploration.
2. Identify the main therapeutic benefits and the active ingredient mentioned for each herbal remedy.
Context:
The text provides an overview of various herbal remedies used for centuries. It mentions that Chamomile contains Bisabolol, which has anti-inflammatory and calming properties. Gingko Biloba, known for its flavonoids and terpenoids, enhances cognitive function and blood circulation. Lastly, Echinacea is recognized for its alkamides, which bolster the immune system.
Reasoning:
1. If a cube’s volume is tripled, by what factor does the length of one of its sides increase?
2. In a two-legged soccer match, Team A wins the first leg at home 3-0, but loses the second leg away 2-5. Who advances to the next round, considering the away goals rule?
math:
1. How do you solve the differential equation $dy/dx+2y=e^{(-2x)}$ given that $y(0)=1$?
2. What is the integral of ($x^{2}+2x+2)/(x^{3}+3x^{2}+3x+1)dx$?
Coding:
1. How can I implement a function in C++ that dynamically allocates a 2D array based on user input sizes, initializes all elements to zero, and then deallocates the memory properly to avoid memory leaks?
2. Write a JavaScript function to fetch data from a given URL, parse the JSON response, and filter the results to return an array of items where a specific key’s value matches a condition.
STEM knowledge:
1. How do you calculate the Schwarzschild radius of a black hole, and what implications does this have for the concept of event horizons in general relativity?
2. Can you explain the process of splicing in eukaryotic gene expression and its significance in the diversity of the proteome?
Humanities/social science knowledge:
1. Discuss the impact of colonial legacies on contemporary political structures in African countries, with examples.
2. Analyze the social and economic consequences of the one-child policy in China.
A.3 Inter-judge Agreement
Figure 11: Cohen’s Kappa Agreement with Majority Vote Before Committee Discussions.
Figure 12: Cohen’s Kappa Agreement with Majority Vote After 1 Round of Committee Discussion.
As shown in Figure 4, the Cohen’s Kappa agreement [29] among judges before committee discussion is very low, averaging 0.10, which indicates slight agreement. We notice that weak model judges and strong model judges has an especially low agreement, such as GPT-4 and Llama -2. This shows that general model capabilities could result in significant performance gaps when used as judges.
A.4 Model selection for the main experiment
Table 7: Model Selection for the Main Experiment. “Newest” and “Strongest” refer to the state at the time of experiments (2024 April). Bolded models are selected for the primary experiment with 7 models. Unbolded models are the ones added during extension.
Model Name Reasons for Inclusion License
GPT-4-0409-Turbo [34] Newest and Strongest in GPT-4 model family Proprietary
Claude-3-Haiku [2] Newest and Cheapest in Claude-3 model family Proprietary
Qwen1.5-72B-chat [3] Strongest in Qwen model family Qianwen LICENSE
Mixtral-8x7b-Instruct-v0.1 [23] Strongest in open-source Mistral small models Apache 2.0
MOE Structure
Yi-34B-Chat [1] Strongest in Yi Model Family Yi License
GPT-3.5-Turbo-0125 [32] Newest in GPT-3.5 Model Family Proprietary
Llama-2-70b-chat [38] Strongest in Llama Model Family Llama 2 Community
GPT-4o-2024-05-13 [33] Newly released model in GPT-4 Model Family Proprietary
Llama-3-70b-chat-hf [30] Newly released model in Llama Model Family Llama 3 Community
Reka-core-20240501 [37] Strongest model in Reka Model Family Proprietary
Command R Plus [15] Strongest model in Command R Model Family CC-BY-NC-4.0
mistral-large-2402 [16] Strongest model in mistral Family Proprietary
deepseek-llm-67b-chat [17] Strongest model in deepseek Family DeepSeek License
Wenxin-5 Famous Chinese model, not on Chatbot Arena Proprietary
GLM-4 Famous Chinese model, not on Chatbot Arena Proprietary
Minimax-abab6.5 Famous Chinese model, not on Chatbot Arena Proprietary
SenseChat-5 Famous Chinese model, not on Chatbot Arena Proprietary
In Table 7, we show all the models selected for the main experiment and expansion. We also include the reasons for selection. Overall, we try to select a representative set of famous models on Chatbot Arena top 20 list. While the Chatbot Arena ranking mostly consists of models with different versions, we only select the strongest or newest model from each model family. Besides the models on Chatbot Arena, we include 4 under-evaluated famous Chinese models to investigate their performances.
A.5 Comparison of baseline methods and Auto-Arena
Table 8: Comparison between Auto-Arena and Other Benchmarks.
Method Manual Construction Freshness Eval. Cost Judge
of Queries per Model
MMLU [22] Yes Static - Answer Accuracy
OpenLLM Leaderboard [6] Yes Static - Answer Accuracy
LC-AlpacaEval [19] Yes Static \$10 Single LLM (GPT-4)
MT-Bench [44] Yes Static \$10 Single LLM (GPT-4)
MT-Bench Hard [26] Yes Frequent Updates \$25 Single LLM (GPT-4)
Chatbot Arena [44] Yes Live Very High Humans
Auto-Arena No Freshly Generated \$5 Committee of LLMs
Table 8 shows a comparison between benchmark evaluation methods and Auto-Arena. Compared to previous methods, the main advantage of Auto-Arena is the zero need for human dataset construction or intervention and the freshness of queries. Another main innovation compared to previous model-based systematic benchmarking procedures is using a committee of LLMs to discuss and vote for a final winner, which introduces diverse viewpoints.
For the evaluation cost, the costs of Auto-Arena are on the same scale as other benchmarks: We note that the primary experiment among 7 models costs around \$35 USD. Therefore, the estimated cost is \$5 per model. As models on the ranking board increase, the costs of conducting debates should grow slowly in log scale, which comes from conducting $nlog_{2}(n)$ pairings when adding 1 model to a ranking of $(n-1)$ models. The evaluation costs, however, shall remain the same as we use a committee of 5 LLMs at all times.